The bolts on the cylinder head of an engine require tightening to a torque of 94 m⋅N .
Part A
If a wrench is 28 cm long, what force perpendicular to the wrench must the mechanic exert at its end?
Express your answer to two significant figures and include the appropriate units.
Part B
If the six-sided bolt head is 15 mm in diameter, estimate the force applied near each of the six points by a socket wrench.
Express your answer to two significant figures and include the appropriate units.
1
Expert's answer
2019-04-01T10:44:24-0400

A.


"F_1=\frac{\tau}{R} =\frac{94}{0.28}=340 N"

B.


"F_2=\frac{\tau}{nr} =\frac{94}{6(0.0075)}=2100 N"